OpenStudy (anonymous):

What number must be added to the polynomial to complete the square? x^2+x

OpenStudy (anonymous):

9

OpenStudy (anonymous):

oh heck no

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

coefficient of x is 1 half of one is 1/2 1/2 squared is 1/4

OpenStudy (anonymous):

1/4

OpenStudy (anonymous):

like a mantra this is...

OpenStudy (anonymous):

half of ommmm is \[\frac{\text{ommmm}}{2}\] \[(\frac{\text{ommmm}}{2})^2\] is \[\frac{\text{ommmm}^2}{4}\]